# Reducing leakage of single-qubit gates for superconducting quantum processors using analytical control pulse envelopes

### v1

@@ Appendix F Single-qubit gate calibration
Here, we describe our procedure for calibrating the parameters of a single-qubit $R_{X}(\pi/2)$-gate implemented with a control pulse of the form $\Omega_{\mathrm{RF}}(t)=\Omega_{\mathrm{I}}(t)\cos(\omega_{\mathrm{d}}t+\varphi)+\Omega_{\mathrm{Q}}(t)\sin(\omega_{\mathrm{d}}t+\varphi)$, where $\Omega_{\mathrm{I}}(t)=A\times\Omega_{\mathrm{I}}^{\mathrm{norm}}(t)$ with $A$ being the in-phase amplitude, $\Omega_{\mathrm{Q}}(t)=-\beta\dot{\Omega}_{I}(t)/\alpha$, and the phase $\varphi$ is either set to zero for DRAG-P or advanced by $\varphi_{z}$ before each gate for DRAG-L. Different calibration experiments are needed for DRAG-P and DRAG-L as DRAG-P tunes $\beta$ to minimize phase errors, whereas DRAG-L calibrates $\beta$ to minimize leakage as discussed in Sec. II.1.
@@
Appendix G Implementation and calibration of control pulse predistortion.
@@
Having determined the decay time $\tau_{1}$, we calibrate the amplitude coefficient $a_{1}$. To achieve this, we again measure the axis shift $\varphi_{\mathrm{s}}$ as a function of the delay $t_{\mathrm{d}}$ between pulses in an I-distortion characterization experiment but this time we predistort the control envelopes using the calibrated decay time $\tau_{1}=8$ ns and try different values of $a_{1}$ as shown in Fig. 8(d). For each value of $a_{1}$, we evaluate a cost function $c(a_{1})=\sum_{i}\varphi_{\mathrm{s}}(a_{1},t_{\mathrm{d},i})^{2}$ based on the axis shift values measured for different delays $\{t_{\mathrm{d},i}\}$. Sweeping $a_{1}$ and trying few different pulse durations, we find that the cost function reaches its minimum at approximately $a_{1}=-0.028$, where its value is an order of magnitude smaller compared to $a_{1}=0$ for pulse durations $t_{\mathrm{p}}\sim 10$ ns. Using the obtained parameters $\tau_{1}=8$ ns and $a_{1}=-0.028$, we illustrate the in-phase envelope of a FAST DRAG-P pulse with and without predistortion in Fig. 8(f).
@@
We point out that the current calibration procedure for the predistortion parameters $\{\tau_{1},a_{1}\}$ is relatively time-consuming, but fortunately the same calibration works across different pulse shapes and gate durations. We observed that the calibrated parameters systematically improved the gate performance over the cooldown, during which the data for the main text was collected. In a subsequent cooldown without any changes to the cabling, the I-distortions measured from the I-distortion characeterization experiment were observed to no longer follow an exponential decay and could not be corrected well using an exponential IIR inverse filter with a single exponential term. After a further thermal cycle without any changes to the cabling, the simple exponential decay was recovered in the I-distortion characterization experiment and essentially the same parameters $\tau_{1}=8$ ns and $a_{1}=-0.028$ were found to provide optimal performance. Further research is needed to understand the exact origins of the distortions, and to develop predistortion methods for correcting distortions beyond the simple exponential filter model (cf. the remaining C-distortions observed in Fig. 5(f)).
Appendix H Further results on leakage error for different pulse shapes
Here, we compare experimental and simulated leakage error as a function of gate duration, see Fig. 9(a), for all the studied pulse shapes: FAST DRAG-L/P, HD DRAG-L/P, Cosine DRAG-L/P, Slepian DRAG-L/P and Gaussian DRAG-L/P. In general, the experimental results agree with the simulations across the different pulse shapes and gate durations. As an exception, the simulated leakage rate for Slepian DRAG-L is lower than measured for the fastest gates as discussed in Appendix I. To obtain the speed limit presented in Fig. 4(a) for each of the pulses, we linearly interpolate the experimental leakage error to estimate the gate duration, for which the leakage error exceeds the threshold value of $L_{\mathrm{g}}=5\times 10^{-5}$. From Fig. 9(b) showing the envelopes of the studied pulses, we see that the envelopes of FAST DRAG-L and HD DRAG-L resemble each other for the used parameters and require a lower peak amplitude compared to Cosine DRAG-L or Gaussian DRAG-L.
@@ Appendix H Further results on leakage error for different pulse shapes
Appendix I Parameter sweeps for FAST DRAG, HD DRAG and Slepian DRAG

### v2

@@ Appendix G Single-qubit gate calibration
Here, we describe our procedure for calibrating the parameters of a single-qubit $R_{X}(\pi/2)$-gate implemented with a control pulse of the form $\Omega_{\mathrm{RF}}(t)=\Omega_{\mathrm{I}}(t)\cos(\omega_{\mathrm{d}}t+\varphi)+\Omega_{\mathrm{Q}}(t)\sin(\omega_{\mathrm{d}}t+\varphi)$, where $\Omega_{\mathrm{I}}(t)=A\times\Omega_{\mathrm{I}}^{\mathrm{norm}}(t)$ with $A$ being the in-phase amplitude, $\Omega_{\mathrm{Q}}(t)=-\beta\dot{\Omega}_{I}(t)/\alpha$, and the phase $\varphi$ is either set to zero for DRAG-P or advanced by $\varphi_{z}$ before each gate for DRAG-L. Different calibration experiments are needed for DRAG-P and DRAG-L as DRAG-P tunes $\beta$ to minimize phase errors, whereas DRAG-L calibrates $\beta$ to minimize leakage as discussed in Sec. II.1. See Fig. 10 for a flow chart of the calibration procedure.
Figure 10: Calibration flow chart. Flow chart of the single-qubit gate calibration. See Appendix H and Fig. 11 for more details on microwave pulse predistortion calibration and Appendix F for more details on hyperparameter selection.
@@
Appendix H Implementation and calibration of control pulse predistortion.
@@
Having determined the decay time $\tau_{1}$, we calibrate the amplitude coefficient $a_{1}$. To achieve this, we again measure the axis shift $\varphi_{\mathrm{s}}$ as a function of the delay $t_{\mathrm{d}}$ between pulses in an I-distortion characterization experiment but this time we predistort the control envelopes using the calibrated decay time $\tau_{1}=8$ ns and try different values of $a_{1}$, as shown in Fig. 11(d). For each value of $a_{1}$, we evaluate a cost function $c(a_{1})=\sum_{i}\varphi_{\mathrm{s}}(a_{1},t_{\mathrm{d},i})^{2}$ based on the axis shift values measured for different delays $\{t_{\mathrm{d},i}\}$. Sweeping $a_{1}$ and trying few different pulse durations, we find that the cost function reaches its minimum at approximately $a_{1}=-0.028$, where its value is an order of magnitude smaller compared to $a_{1}=0$ for pulse durations $t_{\mathrm{p}}\sim 10$ ns. Using the obtained parameters $\tau_{1}=8$ ns and $a_{1}=-0.028$, we illustrate the in-phase envelope of a FAST DRAG-P pulse with and without predistortion in Fig. 11(f).
@@
We point out that the current calibration procedure for the predistortion parameters $\{\tau_{1},a_{1}\}$ is relatively time-consuming, but fortunately the same calibration works across different pulse shapes and gate durations. We observed that the calibrated parameters systematically improved the gate performance over the cooldown, during which the data for the main text was collected. In a subsequent cooldown without any changes to the cabling, the I-distortions measured from the I-distortion characeterization experiment were observed to no longer follow an exponential decay and could not be corrected well using an exponential IIR inverse filter with a single exponential term. After a further thermal cycle without any changes to the cabling, the simple exponential decay was recovered in the I-distortion characterization experiment and essentially the same parameters $\tau_{1}=8$ ns and $a_{1}=-0.028$ were found to provide optimal performance.
We also attempted to correct the C-distortions observed in Fig. 5(f) by assuming the C-distortions to be described by a cross-quadrature exponential IIR filter model of the form
$y(t)=u(t)\times(1+\mathrm{i}\sum_{j}\tilde{a}_{j}\mathrm{e}^{-t/\tilde{\tau}_{j}}),$ (41)
where $y(t)=y_{\mathrm{I}}(t)+\mathrm{i}y_{\mathrm{Q}}(t)$ is the complex envelope seen by the qubit, $u(t)=u_{\mathrm{I}}(t)+\mathrm{i}u_{\mathrm{Q}}(t)$ is the step function in both $I$ and $Q$ components, and $\{\tilde{a}_{j},\tilde{\tau}_{j}\}$ are the distortion parameters. However, we were not able reduce coherent errors observed in the C-distortion characterization experiment by predistorting the control pulses under the distortion model of Eq. (41) despite optimizing the distortion parameters. Hence, a more appropriate model for the C-distortions would be needed to enable successful predistortion since our distortion characterization experiments require the assumption of a specific filter model. Therefore, further research is needed to understand the exact origins of the distortions and to devise more appropriate filter models. One promising distortion model to be tested in the future are control line reflections that could result in both I- and C-distortions due to an additional phase accrued by the reflected pulses. An alternative direction of research is to develop methods for characterizing the full transfer function of the control lines, e.g., by extending the methods of Ref. [75].
Appendix I Further results on leakage error for different pulse shapes
Here, we compare experimental and simulated leakage error as a function of gate duration, see Fig. 12(a), for all the studied pulse shapes: FAST DRAG-L/P, HD DRAG-L/P, Cosine DRAG-L/P, Slepian DRAG-L/P and Gaussian DRAG-L/P. In general, the experimental results agree with the simulations across the different pulse shapes and gate durations. As an exception, the simulated leakage rate for Slepian DRAG-L is lower than measured for the fastest gates as discussed in Appendix J. To obtain the speed limit presented in Fig. 4(a) for each of the pulses, we linearly interpolate the experimental leakage error to estimate the gate duration, for which the leakage error exceeds the threshold value of $L_{\mathrm{g}}=5\times 10^{-5}$. From Fig. 12(b) showing the envelopes of the studied pulses, we see that the envelopes of FAST DRAG-L and HD DRAG-L resemble each other for the used parameters and require a lower peak amplitude compared to Cosine DRAG-L or Gaussian DRAG-L.
@@ Appendix I Further results on leakage error for different pulse shapes
Appendix J Parameter sweeps for FAST DRAG, HD DRAG and Slepian DRAG